INAF queue observing, 2018 November 07-08

Observer: R. Carini, M. Faccini
Telescope Operator: D. Gonzal-Huerta
Support Astronomer: B. Rothberg
Instrument: LUCI1 (LS mode), LUCI2

Summary

Night duration (12\xB0 el twilight) 1:15-12:45 UT = 11h 30m

Observing time: 9h 40min (84%)

Weather time loss: h m (0%)

technical time loss 1h 50m (16%)

Data Summary

At the sunset the sky is clear, wind speed about 10 m/s, humidity about 50%

Project_RUN Objects Instrument Configuration ExpTime /Time_tot Images Notes
Dark   luci1 and luci2        
Pinna LO Peg Luci2 AO N30; FeII, P-beta, Ks, H2 30/30 m luci2.0029-0076 completed but some images have the star elongated
Pinna HD16760 LUCI2 AO N30; FeII, P-beta, Ks, H2 30/30 m luci1.0078-0124 seeing variable and high, to be repeated
Perna telluric HIP16111 luci2 mos G200LoRes; zjspec   luci1.0128-0131  
Perna RCSGA327 luci2 mos G200LoRes; zjspec 1.2/3h luci2.0160-00188 the last 4 images have seeing very high, to be repeated
Nisini telluric HIP17971 luci2 mos G210Hires; J,H,K luci2.0194-0205  
Nisini field 3 luci2 mos J,H,K 1.5/3h luci2.0210-0233  
Nardini telluric HIP50459 luci1 ls G150Ks; Ks   luci1.0019-0022  
Nardini q3 SDSS J0945 luci1 ls G150Ks; Ks 0.67/1.33 h luci1.0028-0043 last image in twilight but should be good
LUCI flats and arc nardini field3, Pinna          

Details

Dark Luci
  • Antoniucci IRAS luci1.20181108.0001-0005, luci2.20181108.0001-0005 dit 60 ndit 1 mer nor
  • luci1.0006-0010, luci2.0006-0010 dit 30 ndit 1 mer nor
  • V1180, Agll354 luci10011-0015,luci2.0011-0015 dit 300 ndit 1 mer nor

00:30 TO opens the dome
  • AO check
  • collimation

01:22 thare are problems to close the loop
  • problem fixed, TO finishes the check
01:51 TO sends the preset to the target LO Peg, Pinna's program
  • 01:58 the AO support closes and opens system, there is issue in the comunication in the AO system
  • 02:02 TO collimates again

02:07 send the script of LO Peg
  • error during the commit
  • 02:09 send again the preset luci2.20181108.0025 test image
  • fwhm 0.54, estimated magnitude 9.1
  • 02:21 error in the script about the subwindow
  • 02:24 MOS error
  • to fix the MOS mask error the loop has to been opened.

03:15 ready to go again
  • TO send again the preset
  • ==luci2.0027-0028==test images
  • fwhm 0.7"
  • put the subwindow
  • 03:33 start science luci2.0029-0076
  • 03:36 fwhm 0.58", seeing on the WFS camera 0.92", estimated magnitude 9.1 mag
  • stars are elongated since imaluci1.0044, after the elongation decreased
  • 03:59, seeing on the DIMM 0.92, on the WFC camera 1.3", fwhm 0.72", estimated magnitude 9.1 mag
  • 04:15 seeing in the DIMM 0.64", on the WFS Camera 1.04", fwhm 0.65", estimated magnitude 9.1 mag
  • 04:23 seeing in the DIMM0.67", on the WFC Camera 0.99", fwhm 0.77", estimated magnitude 9.1 mag

04:28 check AO

04:32 slew to the target HD16760
  • 04:40 error during commit
  • problem with the AO system

05:00 start again the sript
  • luci1.0077 test ima
  • sci luci1.0078-0124
  • 05:08, seeing on the DIMM 0.94", on th WFC camera 0.94", fwhm 0.74", estimated mag 9.0
  • 05:22 seeing on the DIMM 1.13", on the WFC 1.12", fwhm 0.9", estimated mag 9.1
  • seeing is increasing
  • 05:26 seeing on the DIMM 1.4", on the WFC 1.4", fwhm 1.1", estimated mag 9.1
  • 05:33 seeing on the DIMM and on the WFC 1.6", fwhm 1.35", estimated mag 9.2
  • 05:39" DIMM=1.55", WFC= 1.6", fwhm=1.0", estimated mag 9.1
  • 05:42 DIMM= 1.2", WFC= 1.2", fwhm=0.8", estimated mag 9.0
  • 05:55 DIMM=1.35", WFC= fwhm=0.95"
  • 05:57 DIMM=1.7" fwhm =1.3"
  • seeing fwhm variable between 0.7" 1.5"

06:03 slew to Nisini's program, field 3
  • 06:05 MOS error

06:13 issue fixed. the target of Nisini is very high, the seeing now is about 1'-1.0', we slew to Perna's target
  • 06:21 Aladin is stuck, we close LUCI panle and open it again
  • acq telluric zj spec ==luci2.0125-0127
  • seeing on the DIMM 0.9"
  • sci telluric luci1.0128-0131=

06:36 slew to the target
  • acq luci2.0132-0159 the last 24 images should be integrated
  • sci luci2.0160-00183
  • seeing very good, fwhm 0.8", on the dimm 1.0"
  • 07:34 fwhm 0.8'-1.0", dimm 1.0"-1.1" but it jumps to higher values due to the wind.
  • 07:59 dimm 1.36" and decreases

08:14 send again the script of Perna luci2.00184-0188
  • seeing variable 1.0"-1.4" with jumps to 2.0''
  • 08:22 the seeing on the DIMM 1.4"-1.7"

08:29 the seeing is too high, about 1.4"". We slew to Nisini's program
  • telluric
  • acq luci2.0189-0190
  • 08:39 problem with the collimation, the TO corrects the collimation.

08:45 send again the script of the telluric HIP17971
  • acq luci2.0191-0193
  • sci luci2.0194-0205
  • fwhm 1.3"

09:16 send the script for field 3
  • acq luci2.206-0209
  • sci luci2.0210-0233
  • seeing about 1.8"
  • 09:52 fwhm 1.3"
  • 10:01 fwhm 1.2"
  • 10:30 fhwm 0.9"
  • 10:35 fhwm 1.5"
  • 10:56 fwhm variable between 1.1"and 1.7"
  • 11:06 the seeing in decreasing, variable around 1.1"

11:25 since the seeing is pretty good we slew to Nardini's program, we start with the telluric for q3
  • acq luci1.0015-0018
  • sci luci1.0019-0022
  • seeing about 1.2 on the DIMM, fwhm 1.0

11:55 target q3 SDSS J094554
  • acq luci1.0023-0027
  • sci =luci1.0028-0043
  • 12:02 seeing about 1.2 on the DIMM

Calibrations

Nisini, field3 filter H
  • luci2.0234-0238 5 flush
  • luci2.0239-0243 halo 1, flat
  • luci2.0244-0245 argon, arc
  • luci2.0246-0247 xenon, arc

Nisini, field 3, filter J
  • luci2.0248-0252 5 flush
  • luci2.0253-0257 halo 1, flat
  • luci2.0258-0259 argon, arc
  • luci2.0260-0261 xenon, arc

Nisini, field 3, filter K
  • luci2.0262-0266 5 flush
  • luci2.0267-0271 halo 1, flat
  • luci2.0272-0273 argon, arc
  • luci2.0274-0275 xenon, arc
  • luci2.0276-0277 neon, arc

Pinna
  • luci2.0278-0282 flush
  • luci2.0283-0287 flat Fe II, halo1
  • luci2.0288-0292 flush
  • luci2.0293-0297 flat H2, halo1
  • luci2.0298-0302 flush
  • luci2.0303-0307 flat Ks
Topic revision: r14 - 13 Dec 2018, RobertaCarini
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback